OpenStudy (anonymous):

Which of the following is the best estimate for checking the given problem? 3 3 + 3 2 = 8 9

OpenStudy (anonymous):

?

OpenStudy (anonymous):

3 1 + 3 1 = 2 3 3 4 + 3 1 = 8 2 3 + 3 = 3 1 + 3 = 2